OpenStudy (anonymous):

Verify 2xydx+(x^2-y)dy=0 -2x^2y+y^2=1

OpenStudy (anonymous):

welcome 2 os

OpenStudy (anonymous):

thank you, can you please help me with that equation? i really have no clue of what i have to do

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."

OpenStudy (anonymous):

go 2 this site and it will help

OpenStudy (anonymous):

thanks, ill read it and if i have any question can i ask you?

OpenStudy (anonymous):

so should i write it as -2xy/(x^2-y)=dy/dx?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and then how do i get the substitution the example is telling to do next?

OpenStudy (john_es):

If you are not using differential equations, then you can solve the problem with implicit derivation. Take \[F(x,y)=-2x^2y+y^2-1=0\]Then \[dF=\frac{\partial F}{\partial x}dx+\frac{\partial F}{\partial y}dy=0\] \[\frac{dy}{dx}=-\frac{\frac{\partial F}{\partial x}}{\frac{\partial F}{\partial y}}\] And the result follow from the last step.

OpenStudy (anonymous):

So the partial derivative of F(x,y) in x is -4x and in y is -2x^2+2y=0? Then i have \[-(-4x)/(-2x^2+2y)\]?

OpenStudy (anonymous):

Im sorry to be such a mess! :(

OpenStudy (john_es):

You should have, \[\frac{\partial F}{\partial x}=-4xy\]\[\frac{\partial F}{\partial y}=2y-2x^2\]Then \[\frac{dy}{dx}=\frac{4xy}{2y-2x^2}=\frac{2xy}{y-x^2}\] And \[2xydx+(x^2-y)dy=0\]

OpenStudy (anonymous):

Thank you so much, sorry to not understand things so fast

OpenStudy (john_es):

Don't worry, all of us have a long way to learn. ;)

OpenStudy (anonymous):

Can i ask you an extra question of another doubt i have? its not related to this problem

OpenStudy (john_es):

Of course.

OpenStudy (anonymous):

Okay, so i have \[me ^{mx}+2e ^{mx}=0\], how do i clear the m?

OpenStudy (anonymous):

well i know its with ln, but the ln goes outside the \[me ^{mx}\] or inside?

OpenStudy (john_es):

Well, I would do the following, \[(m+2)e^{mx}=0\Rightarrow m=-2\]

OpenStudy (anonymous):

oh boy, how could i not see it? haha now i really made a fool of myself.

OpenStudy (john_es):

Don't worry, remember the exponential could not reach the 0 value, it is an asymptote.

OpenStudy (anonymous):

Okay, thank you so much for your help

OpenStudy (john_es):

You're welcome. One thing more, the link from previous answers (@kathy0514) is not correct, because the equation in your problem is not homogeneus. However the solved equation from the example is homogeneus.

OpenStudy (anonymous):

oh okay, well i didnt get the part that it tells me to find the substitution, so i kind of got stuck in there.

OpenStudy (john_es):

Still, you can solve it by using differential equations. The equation is exact, so you can solve using the method from this video, http://www.youtube.com/watch?v=bwASJWS8ltM Follow the link only if you have taken the lessons of exact differential equations.

OpenStudy (anonymous):

not yet, well see that topic next week, so i guess ill wait a little bit more

OpenStudy (john_es):

Ok, then problem solved ;)

OpenStudy (anonymous):

perfect, thank you for your time! and have a great day!

OpenStudy (john_es):

Same to you, ;)